OpenStudy (anonymous):

The sum of five consecutive integers is 75. What is the third number? Please don't give me the answer right away....

OpenStudy (anonymous):

x+(x+1)+(x+2)+(x+3)+(x+4)=75

OpenStudy (anonymous):

yes...Can you help me figure out the answer?

OpenStudy (anonymous):

Artist, do you see how he arrived at that equation?

OpenStudy (anonymous):

Yes...I understand....

OpenStudy (anonymous):

Ok, so how many x's do you have there on the left side.

OpenStudy (anonymous):

five

OpenStudy (anonymous):

Ok, so 5x. And what is the sum of the non-x's on the left side.

OpenStudy (anonymous):

Protip: Use (x-2) + (x-1) + x + (x+1) + (x+2)

OpenStudy (anonymous):

4....

OpenStudy (anonymous):

4+3+2+1 is not 4.

OpenStudy (anonymous):

sorry...10

OpenStudy (anonymous):

Ok, so the left side is 5x + 10 , and that equals 75. What is x?

OpenStudy (anonymous):

umm...18?

OpenStudy (anonymous):

No. What are the steps to solving \[5x + 10 = 75\]

OpenStudy (anonymous):

What do you do first?

OpenStudy (anonymous):

Umm....multiplication, 5*x?

OpenStudy (anonymous):

No. Isolate the x term, but subtracting 10 from both sides.

OpenStudy (anonymous):

err by subtracting I mean.

OpenStudy (anonymous):

ooh...okay...duh

OpenStudy (anonymous):

Then divide both sides by 5 is the second step. What is x?

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

Ok, so 13 is x is the first number, then what is the third number?

OpenStudy (anonymous):

39...

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

The sum of 5 sequential integers . The first integer is 13, what is the integer that comes after 13?

OpenStudy (anonymous):

Duh,14...so the third number is 15?

OpenStudy (anonymous):

Yes. Exactly.

OpenStudy (anonymous):

13 + 14 + 15 + 16 + 17 = 75

OpenStudy (anonymous):

Thank you...

OpenStudy (anonymous):

Of course! =)

OpenStudy (anonymous):

I fanned you! Can you fan me? you'll be my tenth fan!

OpenStudy (anonymous):

Not that it matters much, but sure ;p

OpenStudy (anonymous):

Thanks...It's kinda neat to have ten fans...

OpenStudy (anonymous):

My method was cleaner *sad face*

OpenStudy (anonymous):

Ooh...sorry....*sympathetic face*

OpenStudy (anonymous):

I agree Newton, but ankur had already started in this vein, and it doesn't matter much either way, so long as the method is understood.

OpenStudy (anonymous):

Yes...I agree...Maybe next time INewton....